# Fakhraddin Alwajih     El Moatez Billah Nagoudi     Gagan Bhatia Abdelrahman Mohamed     Muhammad Abdul-Mageed

### v1

@@ 1 Introduction
Figure 1: Comparison between the performance of Peacock models on SEED-Benchmark dimensions.
@@
To summarize, our contributions in this paper are as follows: (1) We introduce a suite of Arabic MMLLs, dubbed Peacock, capable of instruction following and visual reasoning, in addition to their intriguing dialectal affinity. For developing Peacock, we use existing vision and language models. We also offer a new language model, AraLLaMA, based on LLaMA2-7B  Touvron et al. (2023). (2) We introduce a diverse collection of Arabic translated datasets carefully curated for the training and evaluation of Arabic MLLMs. (3) We adapt the popular LLaVA Liu et al. (2023b) benchmark and SEED-Bench Li et al. (2023d) for Arabic MLLMs evaluation. (4) We present Henna, a benchmark designed to measure model capabilities in interpreting images related to Arabic culture.
@@ 4.1 Translation and Filtering Pipeline
A significant challenge for Arabic MLLMs is lack of available resources, which is due to the difficulty of retrieving relevant Arabic image-text pairs from the internet at scale and absence of suitable image-text relevance filtering methods similar to that of CLIP Radford et al. (2021) [^2] CLIP was used in filtering many English web scraped large-scale datasets Ordonez et al. (2011); Sharma et al. (2018); Changpinyo et al. (2021).. To address this resource gap, we introduce a careful translate-and-filter pipeline for converting publicly available image-text datasets into Arabic without losing data quality. To this end, we adopt the latest version of Google translate API Google Cloud . We follow  Mohamed et al. (2023) in further assuring high quality of acquired translations using a multilingual sentence embedding modelLaBSE Feng et al. (2020). We calculate the similarity of embeddings between the original and translated sentences, keeping translations with a minimum similarity threshold >80absent80>80> 80% to source data. Figure 4 demonstrates the filtering pipeline. Table 1 presents the statistics of the translated data before and after the filtering process. As the table indicates, the reduction ratio due to filtering is approximately 20%. We provide details about our translation method in Appendix A.3, and sample translations illustrating variations in quality ranging from good to moderate to poor in Figure A.1 (also in Appendix A.3).
Dataset        Type Non Filtered Filtered
Train Dev Test Train Dev Test
COCO2017 Captioning 590k 25k 527k 22k
LLaVA Pretrain (LCS) Captioning 558k 389k
LLaVA instruct 150k Instructions 271k 204k
VQAv2 VQA 440k 214k 351k 172k
OKVQA VQA 9k 4k 7k 3k
GQA VQA 938k 132k 638k 89k
SEED-Bench Benchmark 14k 7k
LLaVA-Bench Benchmark 90 90
Henna-Bench Benchmark 1k
Table 1: Breakdown of our publicly released pretraining and instruction finetuning Arabic datasets. Only the filtered datasets highlighted in green and cyan are used for pretraining and instruction finetuning, respectively.
@@ 5.2.1 General VQA
Non Filtered Filtered Non Filtered Filtered Non Filtered Filtered
Baseline mBlip mT0-XL-5B 38.55 8.59 18.18 35.95 50.45
@@
In general VQA tasks, the challenge involves answering textual questions about images, requiring learning and integrating visual and textual information. This demonstrates a deep understanding of the interconnectedness between the two modalities. To evaluate performance in general VQA, our validation process includes three datasets: VQAv2 Goyal et al. (2017), OKVQA Marino et al. (2019), and GQA Hudson and Manning (2019). Notably, evaluation of English VQA tasks is typically performed through online platforms by submitting results. However, this option is unavailable for Arabic-translated data because these platforms only support English and the original datasets. Since the test sets do not contain ground truth labels, we evaluate held-out validation sets. We follow Geigle et al. (2023) in using “exact match accuracy with open generation” to evaluate our models’ output. The metric considers an answer correct if it matches any of the ground truth answers. Table 2 shows model accuracy on these datasets under the zero-shot setting for both the filtered and unfiltered versions of the datasets.
@@
Furthermore, we observe that the choice of the underlying LLM is has a significant impact on performance. This is the case if we compare integrating AraLLaMA to AceGPT in our overall MLLMs. Specifically, the InstructBlip model integrated with AraLLaMA demonstrates superior performance across all tasks and datasets when using either filtered or unfiltered data in training. This performance disparity is likely attributable to the inherent differences in how these LLMs handle Arabic, with AraLLaMA being more effective due to its extensive training and its ability to align with visual information. In addition, it is worth noting that the performance of Peacock models varies considerably depending on the task, with a general trend of models performing better on the VQAv2 task than on OKVQA and GQA. Such variations can be attributed to each task’s inherent complexities and specific requirements, including the sophistication of the presented questions and the nature of the required visual understanding.
@@ 5.2.3 SEED Benchmark
Table 4: Evaluation of Peacock models on SEED-Bench across various attributes: Instance Attributes (IA), Instance Identity (II), Instance Interaction (IN), Instance Location (IL), Instances Counting (IC), Scene Understanding (SU), Spatial Relation (SR), and Visual Reasoning (VR).
For our third benchmark, we adapt the SEED-Bench for Arabic and use it to evaluate our models. Table 4 and Figure 1 present an evaluation of Peacock models across a broad spectrum of visual understanding dimensions within SEED-Bench, where a diverse range of performance efficiencies is observed. LLaVA-AraLLaMA emerges as a particularly robust model, excelling in visual reasoning and scene understanding with accuracy scores of 68.88% and 65.52%, respectively. However, it displays weaknesses in spatial relations and instance location. Mirroring this trend, LLaVA-AceGPT showcases strengths in scene understanding and visual reasoning (66.72% and 71.6%, respectively), but marginally underperforms in instance interaction and spatial relations compared to LLaVA-AraLLaMA. In contrast, InstructBlip-AraLLaMA, while proficient in scene understanding and Visual Reasoning (59.34% and 64.95%), falls short in Instance Attributes and Counting, resulting in a lower overall accuracy of 46.43%. InstructBlip-AceGPT, the model with the most modest performance, achieves its best results in visual reasoning and instance interaction (51.06% and 43.3%), but struggles significantly with instance counting and scene understanding.
@@ 5.2.4 Henna Benchmark
Figure 5: Examples of responses from Peacock and GPT-V4 regarding an image related to Yemeni culture.
@@ 5.2.5 Case Study with Egyptian Dialect
Figure 7: Human evaluation results on the accuracy and authenticity of model responses to questions about images in Egyptian dialect. ”IB-AraLLaMA ” denotes our InstructBlip-AraLLaMA model, and ”IB-AceGPT” refers to our InstructBlip-AceGPT model. ”GPT-4V DI” is the GPT-4V model explicitly instructed to respond in the Egyptian dialect.”GPT-4V” represents the GPT-4V model, which is given a question in the Egyptian dialect, similar to how Peacock models are instructed.
@@ 7 Limitations
  • Additionally, translation errors can significantly impact the model’s performance and propagate through the training data. We have identified several such errors in the model’s responses. For example, the English word ’sitting’ typically indicates the location of an object. However, the Google API often mistranslates it to suggest that the object is lying down, as seen in the translation of ’The train is sitting at the station’ to \<القطار جالس في المحطة¿, where \<جالس¿ inaccurately implies that the train is lying down.
@@ A.3 Translation and Filtering
We also assessed translations by the No Language Left Behind (NLLB) project Costa-jussà et al. (2022) In the literature, ElJundi et al. (2020) found that Google translates API translations to be sub-optimal, while Mohamed et al. (2023) showed that NLLB performs better than the free Google translate API integrated into Google Sheets. However, our investigation reveals that the latest Google Translate API has superior performance and provides higher-quality translations than NLLB. This observation aligns with the recent findings of  Kadaoui et al. (2023).
@@ A.6 Henna Generation Pipeline and Examples
We selected an image for each attraction and used GPT-4V to generate ten questions. Each image is accompanied by its Wikipedia article to enrich the context with comprehensive descriptions when prompting GPT-4V for generation. This approach yielded a minimum of ten images per country, resulting in a total of 1,132 question-answer pairs across all countries. An example of the dataset generation process is illustrated in Figure A.3 and Figure A.4 demonstrates four randomly selected images with a generated pair of a question and their answers.
@@ A.7 Qualitative Analysis
Figure A.7: InstructBlip-AraLLaMA Examples.
Figure A.8: InstructBlip-AceGPT Examples.
Figure A.9: LLaVA-AraLLaMA Examples.

### v2

@@ 1 Introduction
Figure 1: Comparison between the performance of Peacock and mBlip models on SEED-Benchmark dimensions.
@@
To summarize, our contributions in this paper are as follows: (1) We introduce a suite of Arabic MLLMs, dubbed Peacock, capable of instruction following and visual reasoning, in addition to their intriguing dialectal affinity. For developing Peacock, we use existing vision and language models. We also offer a new language model, AraLLaMA, based on LLaMA2-7B  Touvron et al. (2023). (2) We introduce a diverse collection of Arabic translated datasets carefully curated for the training and evaluation of Arabic MLLMs. (3) We adapt the popular LLaVA Liu et al. (2023b) benchmark and SEED-Bench Li et al. (2023d) for Arabic MLLMs evaluation. (4) We present Henna, a benchmark designed to measure model capabilities in interpreting images related to Arabic culture.
@@ 4.1 Translation and Filtering Pipeline
A significant challenge for Arabic MLLMs is lack of available resources, which is due to the difficulty of retrieving relevant Arabic image-text pairs from the internet at scale and absence of suitable image-text relevance filtering methods similar to that of CLIP Radford et al. (2021) [^2] CLIP was used in filtering many English web scraped large-scale datasets Ordonez et al. (2011); Sharma et al. (2018); Changpinyo et al. (2021).. To address this resource gap, we introduce a careful translate-and-filter pipeline for converting publicly available image-text datasets into Arabic without losing data quality. To this end, we adopt the latest version of Google translate API Google Cloud . We follow  Mohamed et al. (2023) in further assuring high quality of acquired translations using a multilingual sentence embedding model LaBSE Feng et al. (2020). We calculate the similarity of embeddings between the original and translated sentences (questions and answers), retaining translations that meet a minimum similarity threshold of greater than 80% for both the question and the answer. Figure 4 demonstrates the filtering pipeline. We provide details about our datasets and translation method in Appendix A.1, and sample translations illustrating variations in quality ranging from good to moderate to poor in Figure A.1 (also in Appendix A.1).
@@ 5.2.1 General VQA
All Filtered All Filtered All Filtered
Baseline mBlip mT0-XL-5B 38.55 50.8 8.59 18.18 35.95 50.45
@@
In general VQA tasks, the challenge involves answering textual questions about images, requiring learning and integrating visual and textual information. This demonstrates a deep understanding of the interconnectedness between the two modalities. To evaluate performance in general VQA, our validation process includes three datasets: VQAv2 Goyal et al. (2017), OKVQA Marino et al. (2019), and GQA Hudson and Manning (2019). Notably, evaluation of English VQA tasks is typically performed through online platforms by submitting results. However, this option is unavailable for Arabic-translated data because these platforms only support English and the original datasets. Since the test sets do not contain ground truth labels, we evaluate held-out validation sets. We follow Geigle et al. (2023) in using “exact match accuracy with open generation" to evaluate our models’ output. The metric considers an answer correct if it matches any of the ground truth answers. Table 1 shows model accuracy on these datasets under the zero-shot setting for both the filtered and unfiltered(All) versions of the datasets.
@@
Furthermore, we observe that the choice of the underlying LLM is has a significant impact on performance. This is the case if we compare integrating AraLLaMA to AceGPT in our overall MLLMs. Specifically, the InstructBlip model integrated with AraLLaMA demonstrates superior performance across all tasks and datasets when using either filtered or unfiltered(All) data in training. This performance disparity is likely attributable to the inherent differences in how these LLMs handle Arabic, with AraLLaMA being more effective due to its extensive training and its ability to align with visual information. In addition, it is worth noting that the performance of Peacock models varies considerably depending on the task, with a general trend of models performing better on the VQAv2 task than on OKVQA and GQA. Such variations can be attributed to each task’s inherent complexities and specific requirements, including the sophistication of the presented questions and the nature of the required visual understanding.
@@ 5.2.3 SEED Benchmark
Table 3: Evaluation of mBlip and Peacock models on SEED-Bench across various attributes: Instance Attributes (IA), Instance Identity (II), Instance Interaction (IN), Instance Location (IL), Instances Counting (IC), Scene Understanding (SU), Spatial Relation (SR), and Visual Reasoning (VR).
For our third benchmark, we adapt the SEED-Bench for Arabic and use it to evaluate our models. Table 3 and Figure 1 present an evaluation of Peacock models across a broad spectrum of visual understanding dimensions within SEED-Bench, where a diverse range of performance efficiencies is observed. LLaVA-AraLLaMA emerges as a particularly robust model, excelling in visual reasoning and scene understanding with accuracy scores of 68.88% and 65.52%, respectively. However, it displays weaknesses in spatial relations and instance location. Mirroring this trend, LLaVA-AceGPT showcases strengths in scene understanding and visual reasoning (66.72% and 71.6%, respectively), but marginally underperforms in instance interaction and spatial relations compared to LLaVA-AraLLaMA. In contrast, InstructBlip-AraLLaMA, while proficient in scene understanding and Visual Reasoning (59.34% and 64.95%), falls short in Instance Attributes and Counting, resulting in a lower overall accuracy of 46.43%. InstructBlip-AceGPT, the model with the most modest performance, achieves its best results in visual reasoning and instance interaction (51.06% and 43.3%), but struggles significantly with instance counting and scene understanding. In contrast to mBlip, which outperforms Peacock models only in one dimension (instance interaction) and achieves the same score in the spatial relation dimension as InstructBlip-AraLLaMA.
@@ 5.2.4 Henna Benchmark
Figure 5: Examples of responses from Peacock and GPT-4V regarding an image related to Yemeni culture.
@@ 5.2.5 Case Study with Egyptian Dialect
Figure 7: Human evaluation results on the accuracy and authenticity of model responses to questions about images in Egyptian dialect. "IB-AraLLaMA " denotes our InstructBlip-AraLLaMA model, and "IB-AceGPT" refers to our InstructBlip-AceGPT model. "GPT-4V DI" is the GPT-4V model explicitly instructed to respond in the Egyptian dialect."GPT-4V" represents the GPT-4V model, which is given a question in the Egyptian dialect, similar to how Peacock models are instructed.
@@ 7 Limitations
  • Additionally, translation errors can significantly impact the model’s performance and propagate through the training data. We have identified several such errors in the model’s responses. For example, the English word ’sitting’ typically indicates the location of an object. However, the Google API often mistranslates it to suggest that the object is lying down, as seen in the translation of ’The train is sitting at the station’ to \<القطار جالس في المحطة>, where \<جالس> inaccurately implies that the train is lying down.
@@ A.3 Translation and Filtering
Dataset        Type All Filtered
Train Dev Test Train Dev Test
COCO2017 Captioning 590k 25k 527k 22k
LLaVA Pretrain (LCS) Captioning 558k 389k
LLaVA instruct 150k Instructions 271k 204k
VQAv2 VQA 440k 214k 351k 172k
OKVQA VQA 9k 4k 7k 3k
GQA VQA 938k 132k 638k 89k
SEED-Bench Benchmark 14k 7k
LLaVA-Bench Benchmark 90 90
Henna-Bench Benchmark 1k
Table A.1: Breakdown of our publicly released pretraining and instruction finetuning Arabic datasets. Only the filtered datasets highlighted in green and cyan are used for pretraining and instruction finetuning, respectively.
@@
In the literature, ElJundi et al. (2020) found that Google translates API translations to be sub-optimal, while Mohamed et al. (2023) showed that No Language Left Behind (NLLB) project Costa-jussà et al. (2022) performs better than the free Google translate API integrated into Google Sheets. However, our investigation reveals that the latest Google Translate API has superior performance and provides higher-quality translations than NLLB. This observation aligns with the recent findings of  Kadaoui et al. (2023).
To evaluate the effectiveness of filtering in our study, we randomly selected a subset of 250 samples from the VQAv2 dataset and performed human annotations to assess the accuracy of question-and-answer (Q&A) translations in both filtered and unfiltered (All) segments. The analysis revealed that out of the 250 samples in the unfiltered segment, 35 were incorrectly translated. In contrast, within the filtered segment, which comprised 207 samples after the application of a filtering mechanism, only 7 were incorrectly translated. The filter effectively removed 43 samples, of which 31 were inaccurately translated, indicating that the majority of removed items were indeed problematic, while 12 were accurately translated.
Precision Recall F1 Score Accuracy
72.09% 88.57% 79.49% 93.52%
Table A.2: Performance metrics for the filtering process
To quantify the effectiveness of the filtering process, we calculated several metrics as shown in Table A.2. These results, presented in Table A.2, demonstrate that the filtering approach substantially enhances the quality of the dataset by effectively identifying and removing incorrectly translated Q&A pairs.
Table A.1 presents the statistics of the translated data before and after the filtering process. As the table indicates, the reduction ratio due to filtering is approximately 20%
@@ A.6 Henna Generation Pipeline and Examples
We selected an image for each attraction and used GPT-4V to generate ten questions. Each image is accompanied by its Wikipedia article to enrich the context with comprehensive descriptions when prompting GPT-4V for generation. This approach yielded a minimum of ten images per country, resulting in a total of 1,132 question-answer pairs across all countries. An example of the dataset generation process is illustrated in Figure A.3 and Figure A.4 demonstrates four randomly selected images with a generated pair of a question and their answers.Moreover, Figure 5(b) shows examples of questions and answers generated by the Henna pipeline from the image depicted in Figure 5(a). We translated these questions and answers into English as a qualitative example and to provide a better assessment of the dataset’s quality.
@@
(a) Sample image from Henna for the Giza Pyramids in Egypt.
(b) Ten questions and answers generated using the image from 5(a) with Henna generation pipeline. All questions and answers have been translated into English to demonstrate the quality of the generated data.